# ALLO: A Photorealistic Dataset and Data Generation Pipeline for

### v1

@@ IV-C Results
The results of all algorithms on the ALLO test set are shown in Table III. Reverse Distillation (Rev. Dist.) achieved the highest pixel AP at 48.4%, followed by STFPM at 32.4% and FastFlow at 31.7%. Compared to its performance on the MVTec dataset, Reverse Distillation’s image AUROC score decreased by 40.4%, and its pixel AUROC dropped by 46.0% on the ALLO dataset. While Reverse Distillation had the highest pixel AP score, it had a fairly low pixel AUROC score indicating a high false positive rate. FastFlow had the highest image and pixel AUROC scores at 66.9% and 86.0% respectively. Generally, the student-teacher methods (STFPM and Rev. Dist.) performed best, followed by the normalizing flow methods (FastFlow and UFlow), while the semi-supervised methods (DRAEM and DSR) performed worse. Example inferences are shown in Fig. 6.
@@ IV-D Ablation Studies
While all algorithms struggled to generalize to the space domain, we improved the performance of most algorithms using dataset-specific tuning, described in Section IV-A. The effect of data augmentation and custom normalization on FastFlow (the algorithm with the highest pixel AUROC) is shown in Table II. FastFlow was able to achieve higher pixel AUROC and pixel AP scores when using custom image statistics than with ImageNet-based normalization values. This reflects how different the scenes in the ALLO dataset are to those in ImageNet and other datasets that use ImageNet normalization values.
@@
Dataset Model FastFlow [33] UFlow [26] CFA [22] DRAEM [23] DSR [40] Rev. Dist. [21] STFPM [38]
MVTec Image AUROC $\uparrow$ 99.4 98.9 99.5 98.0 98.2 98.5 95.5
Pixel AUROC $\uparrow$ 98.5 98.7 98.5 97.3 - 97.8 97.0
Pixel AP $\uparrow$ - - - 68.4 70.2 - -

### v3

@@ IV-C Results
The results of all algorithms on the ALLO test set are shown in Table III. FastFlow achieved the highest pixel AP at 31.9%, followed by UFlow at 28.4% and DSR at 26.9%. Compared to its performance on the MVTec dataset, FastFlow’s image AUROC score decreased by 32.5%, and its pixel AUROC dropped by 12.5% on the ALLO dataset. FastFlow also had the highest image and pixel AUROC scores at 66.9% and 86.0% respectively. Generally, the normalizing flow methods (FastFlow and UFlow) performed best, followed by the student-teacher methods (STFPM and Rev. Dist.), while the semi-supervised methods (DRAEM and DSR) performed worse. Example inferences are shown in Fig. 6.
@@ IV-D Ablation Studies
While all algorithms struggled to generalize to the space domain, we improved the performance of most algorithms using dataset-specific tuning, described in Section IV-A. The effect of data augmentation and custom normalization on FastFlow (the algorithm with the highest pixel AP) is shown in Table II. FastFlow was able to achieve higher pixel AUROC and pixel AP scores when using custom image statistics than with ImageNet-based normalization values. This reflects how different the scenes in the ALLO dataset are to those in ImageNet and other datasets that use ImageNet normalization values.
@@
Dataset Model FastFlow [33] UFlow [26] CFA [22] DRAEM [23] DSR [40] Rev. Dist. [21] STFPM [38] SuperSimpleNet [47]
MVTec I.AUROC $\uparrow$ 99.4 98.9 99.5 98.0 98.2 98.5 95.5 98.0
P.AUROC $\uparrow$ 98.5 98.7 98.5 97.3 - 97.8 97.0 -
P.AP $\uparrow$ - - - 68.4 70.2 - - -